A bus and car travel in opposite directioms. After 3 hours, they were 300miles apart. The car is traveling 15mph faster than the bus. How fast was the bus and car traveling?

Answer by ptfile(81)

Let 3hrs = the duration of the vehicles traveled
x+15 = the car's speed
x = the bus's speed
300mi= total distance covered
A bus and car travel in opposite directions. After 3 hours, they were 300miles apart. The car is traveling 15mph faster than the bus.
Distance = (rate)(time)
(x+15)(3)+ (x)3 = 300
3x+45 + 3x = 300
6x = 255
x = 42.5
Substitute x with 42.5
42.5+15 = 57.5
The car's speed is 57.5mph while the bus's speed is 42.5mph.
Check
3 (57.5) + 3 (42.5) = 300
172.5 + 127.5 = 300
300 = 300